OpenStudy (anonymous):

the number N, in millions, of hearing impaired individuals of age x can be approximated by N=-0.00006x^3+0.006x^2-0.1x+1.5. Approximate to the nearest million the number of hearing impaired individuals of age 25 and 45

OpenStudy (kropot72):

Have you tried to calculate the answers?

OpenStudy (anonymous):

no looking at my notes it is a scientific notation but thats all I know

OpenStudy (kropot72):

Can you try calculating the first term (-0.00006x^3) by sustituting 25 for x?

OpenStudy (anonymous):

0.9375

OpenStudy (anonymous):

or 0.938

OpenStudy (kropot72):

\[-0.00006x ^{3}=-(6\times 10^{-5})\times x ^{3}\] \[-(6\times 10^{-5})\times 25^{3}=-0.9375\]

OpenStudy (anonymous):

Ok I got a final answer of 3.69 .....would the nearest million be 4?

OpenStudy (anonymous):

wait the first number is neg so Im wrong

OpenStudy (kropot72):

The second term can be written as: \[6\times10^{-3}\times x ^{3}\] Can you substitute 25 for x and calculate?

OpenStudy (anonymous):

3.75

OpenStudy (kropot72):

Good work. So the calculation for age 25 years is: -0.9375 + 3.75 -2.5 + 1.5

OpenStudy (anonymous):

1.8125

OpenStudy (anonymous):

would I round it to 2 million?

OpenStudy (anonymous):

or 1.81

OpenStudy (kropot72):

Gr8! Yes, you are required to round to 2 million. Do you want to try for age 45 and post back the answer for me to check?

OpenStudy (anonymous):

awesome THANK YOU!!!! Sure give me a couple mins

OpenStudy (anonymous):

15 million?

OpenStudy (kropot72):

Check against my result for each term as follows: -5.4675 + 12.15 - 4.5 + 1.5

OpenStudy (anonymous):

yes that what I got answer 14.6175

OpenStudy (kropot72):

Did you get all the signs right? I get 3.6825 million which rounds to 4 million.

OpenStudy (anonymous):

I recalculated and got the same, I supose I hit the wrong sign somewhere

OpenStudy (anonymous):

same as in 3.6825

OpenStudy (kropot72):

Wonderful. Thanks for perservering :)

OpenStudy (anonymous):

Auh THANK YOU for walking me through it!! Your a life saver :). This is my last class to get my associates degree and I am HORRIBLE with math

OpenStudy (kropot72):

Good luck with your studies. I had to work very hard to get my math grades up so I know how you feel

OpenStudy (anonymous):

thanks